OpenStudy (anonymous):

Solve for w in P = 2w + 2l, if P = 38 and l = 12.

OpenStudy (anonymous):

lmao im just too lazy to do math ..

OpenStudy (.cheesecake.):

38=2w+2(12) 2 x 12=?

OpenStudy (koikkara):

I would solve this way.. P = 2w + 2l P = 38 l = 12 38 = 2w + 2 * 12 38 = 2w + 24 14 = 2w 7 = w

OpenStudy (.cheesecake.):

@Koikkara can you stop telling the user answers

OpenStudy (anonymous):

no i aprecciate it lol

OpenStudy (koikkara):

@akroayee Learn to solve like this.... they just get jealous of me... lol, Have a nice day !

OpenStudy (koikkara):

@akroayee Next 3 questions, your turn... bcuz i know what you gonna ask.

OpenStudy (welshfella):

@akroayee I hope you study the solution so you are able to do a similar prolem on your own. Thats more important than just knowing the answer.

pooja195 (pooja195):

It's not jealousy, we have rules on this site. Follow them or simply get banned. "Give Help, Not Answers-I will encourage and guide those needing help, and not just give them an answer" Don't make a post in openstudy feedback asking why you got suspended when you do.

OpenStudy (anonymous):

you guys suck

OpenStudy (anonymous):

ive got another question .-.

OpenStudy (.cheesecake.):

Feel free to make a new post and post it :)

OpenStudy (welshfella):

You attempt to help people and they make remarks lkie that!!

OpenStudy (anonymous):

yes